# Near-Field Localization and Phase Shift Optimization for RIS-Assisted Non-Ideal OFDM Systems

### v2

@@ II System Model
    where $\bm{h}$ and $\bm{v}$ represent the LoS part of the channel and the combined noise, respectively. $\bm{h}$ contains the known part of the channel from AN to RIS and the unknown part of the channel from RIS to UE. Each item of it can be expressed by $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path loss from AN to RIS, and from RIS to UE, respectively. $\bm{w}[r]$ is the phase shift of the $r$-th element of RIS, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ represents the delay from AN to the $r$-th element of RIS, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th element of RIS to UE, with $c$ representing the light speed,
  • $\bm{v}\triangleq\left[\bm{v}[0],\bm{v}[1],\cdots,\bm{v}[N-1]\right]$ is the combined additive noise at UE, and follows the distribution $\bm{v}\sim\mathcal{CN}(\bm{0}_{N\times 1},\sigma^{2}\mathbf{I}_{N})$.
Drawing upon the attributes of PN observed in practical oscillators, it can be effectively modeled as a Wiener process [22], i.e.,
$\bm{\theta}[n]=\bm{\theta}[n-1]+\bm{\Delta}[n],$ (5)
where $\bm{\Delta}[n]$ is a real Gaussian variable with a distribution of $\bm{\Delta}[n]\sim\mathcal{N}(0,\sigma_{\Delta}^{2})$. As in [21], it is assumed that $\bm{\theta}[-1]=0$, and then we can obtain that the PN vector $\bm{\theta}\triangleq\left[\bm{\theta}[0],\bm{\theta}[1],\cdots,\bm{\theta}[N-1]\right]^{\top}$ follows a distribution of $\bm{\theta}\sim\mathcal{N}(\bm{0}_{N\times 1},\bm{\Psi})$, where $\mathbf{\Psi}$ can be expressed by
$\mathbf{\Psi}=\sigma_{\Delta}^{2}\begin{bmatrix}1&1&1&\cdots&1&1\\ 1&2&2&\cdots&2&2\\ 1&2&3&3&\cdots&3\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&2&3&\cdots&N-1&N\end{bmatrix}.$ (6)
III Proposed Joint Position, CFO and PN Estimation
In previous works [12, 16], localization was based on frequency-domain signals. However, in our problem, the introduction of CFO and PN would result in ICI if the signal is transformed into the frequency domain, which will increase the difficulty of our estimation. Therefore, in this paper, we conduct the estimation in the time domain. In this section, we propose an iterative algorithm based on MAP criterion and GD algorithm, which jointly estimates the normalized CFO, PN, and position of UE.
By defining parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, whereas $\bm{\theta}$ is a random variable, and the others are deterministic variables, the posterior distribution of $\bm{\lambda}$ given the received signal $\bm{y}$ can be written as
@@ III Proposed Joint Position, CFO and PN Estimation
After removing the constant terms, the object function can be expressed as $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, and our goal is to achieve the joint estimation of $\bm{\lambda}$ by optimizing the following unconstrained function.
$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}).$ (8)
@@ III-A1 Approximation of PN vector
Figure 2: Eigenvalues of $\bm{\Psi}$ at $N=32$ and $\sigma_{\Delta}^{2}=10^{-3}$.
From (4), it can be observed that there are multiple parameters that need to be jointly estimated. In order to alleviate the computational complexity at UE, we take advantage of the properties of the PN vector to approximate it. This approach results in a reduction of unknown parameters that require estimation. As depicted in Fig. 2, the eigenvalues of $\bm{\Psi}$ mostly cluster around zero. This observation allows us to approximate the PN vector as follows
@@
where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN parameters, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transformation matrix. To derive $\bm{\Pi}$, we first perform singular value decomposition (SVD) on $\bm{\Psi}$, which is given by $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$, and $\bm{e}$ represents the eigenvalues of $\bm{\Psi}$ arranged in descending order. Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$. In the subsequent analysis, $\bm{\eta}$ is estimated instead of $\bm{\theta}$.
@@ III-A2 Estimation of approximate PN vector
After approximating the PN parameter $\bm{\theta}$, the objective function can be rewritten as
@@
where $\bm{\mu}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}$ is the received signal at UE without additive noise, and the noise variance $\sigma^{2}$ is assumed to be known.
Due to the typically small PN variance in practical oscillators, the elements of $\bm{\Lambda}_{\bm{\theta}}$ can be approximated by employing a first-order Taylor series expansion, resulting in $e^{j\bm{\theta}[n]}\approx 1+j\bm{\theta}[n]$. Consequently, we can approximate $\bm{\Lambda}_{\bm{\theta}}$ as $\bm{\Lambda}_{\bm{\theta}}\approx\mathbf{I}_{N}+j{\rm Diag}(\bm{\Pi\eta})$. With this approximation, the objective function in (III-A2) can be reformulated as:
@@
where $\bar{\bm{y}}\triangleq\bm{y}-\sqrt{P}\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$, and $\mathbf{Q}\triangleq j\sqrt{P}{\rm Diag}(\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]})\bm{\Pi}$. The superscript $q$ represents the round number of iterations.
Subsequently, by setting the gradient of (III-A2) equal to zero, i.e.,
@@ III-B CFO Estimation
To estimate the CFO, we solve the following unconstrained problem
@@
In order to enhance the tractability of the problem presented in (14) and obtain a closed-form solution, we employ a Taylor series approximation once more. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
@@
where $\bm{d}\triangleq\sqrt{P}\hat{\bm{\Lambda}}_{\bm{\theta}}^{[q+1]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$. Upon setting the gradient of (17) equal to zero, and solving for $\phi$, we can obtain the closed-form expression of $\hat{\phi}$ at the $(q+1)$-th iteration as
@@ III-C Position Estimation
For position estimation, we employ the GD algorithm. Due to the similarity in the derivative expressions of the objective function with respect to three positional parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them. The expression for the partial derivative of the objective function with respect to $\xi$ is as follows
@@
where each term of $\frac{\partial\bm{h}}{\partial\xi}$ can be calculated as follows
@@
$\left.\quad+(d_{Ru}cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
The general iterative joint estimation algorithm is outlined in Algorithm 1.
Algorithm 1 Iterative joint estimation algorithm.
1:  Optimize RIS phase shift $\bm{w}$ using the method to be introduced in Section V.
2:  Initialize estimation of normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, step length $k_{p}$, convergence accuracy for outer loop $\epsilon_{o}$, convergence accuracy for inner loop $\epsilon_{i}$, and calculate corresponding polar coordinate of $\hat{\bm{p}}_{u}$ using (1).
3:  repeat
4:     Update $\hat{\bm{\eta}}$ with $\hat{\phi}$, $\hat{\bm{p}}_{u}$ being fixed using (13);
5:     Update $\hat{\bm{\theta}}$ using (9);
6:     Update $\hat{\phi}$ with $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$ being fixed using (18);
7:     repeat
8:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
9:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
10:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
11:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration.
12:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration. $\epsilon_{i}$ and $\epsilon_{o}$ are both pre-set convergence accuracy and should vary with changes in transmit power.
13:  Calculate Cartesian coordinate of $\hat{\bm{p}}_{u}$ using (2).
III-D Estimation Performance Evaluation
For position estimation, we use root mean square error (RMSE) as the performance metric. For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
@@ III-D Estimation Performance Evaluation
where $M_{e}$ is the Monte Carlo experiment amount, and $\hat{\bm{p}}_{u,m}$ represents the estimation of UE position in the $m$-th trail.
For CFO and PN estimation, it can be straightforwardly observed from (4) that, the received signal $\bm{y}$ is not altered under a common phase rotation $\epsilon$, i.e.,
$\hat{\phi}\rightarrow\phi-\epsilon,\hat{\bm{\theta}}\rightarrow\bm{\theta}+\bm{\epsilon},$ (25)
where $\bm{\epsilon}[n]=\frac{2\pi n\epsilon}{N}$. Equation (25) indicates that there exists ambiguity in estimating CFO and PN. The ambiguity leads to difficulties in achieving high estimation accuracy for individual parameters of CFO and PN. However, in this paper, we are not focusing on the separate estimation accuracy of CFO and PN but rather on the estimation accuracy of UE position. Therefore, it is acceptable as long as the overall estimation accuracy of CFO and PN is reduced. In this paper, to evaluate the overall estimation accuracy of CFO and PN, we utilize a novel joint mean square error (MSE) measure, which can be calculated as
@@
where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$.

### v3

@@ II System Model
where $\bm{v}\triangleq\bm{\mu}_{NLoS}+\tilde{\bm{v}}$ is Gaussian distributed, i.e, $\bm{v}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N})$, with $\sigma^{2}=\sigma_{AWGN}^{2}+\sigma_{NLoS}^{2}$.
Each element of $\bm{h}$ can be written as $\bm{h}[n]=\sum_{r=1}^{N_{R}}\sqrt{\rho_{aR,n,r}}\sqrt{\rho_{Ru,n,r}}\bm{w}[r]e^{-j2\pi\frac{nB}{N}(\tau_{ar}+\tau_{ru})}$, where $\sqrt{\rho_{aR,n,r}}=\frac{\sqrt{G_{a}}\lambda_{n}}{4\pi d_{ar}}$ and $\sqrt{\rho_{Ru,n,r}}=\frac{\sqrt{G_{t}}\lambda_{n}}{4\pi d_{ru}}$ are the free-space path losses from the AN to the $r$-th RIS element, and from the $r$-th RIS element to the UE, respectively. $\bm{w}[r]$ is the phase shift of the $r$-th RIS element, $\tau_{ar}=\frac{d_{ar}}{c}=\frac{\|\bm{p}_{r}-\bm{p}_{a}\|}{c}$ is the delay from the AN to the $r$-th RIS element, while $\tau_{ru}=\frac{d_{ru}}{c}=\frac{\|\bm{p}_{u}-\bm{p}_{r}\|}{c}$ denotes the delay from the $r$-th RIS element to the UE, with the light speed $c$.
III Proposed Joint CFO, PN and UE Position Estimation Algorithm
In previous works [27, 30], localization was based on frequency-domain signals, and different subcarriers’ signals were often assumed to be independent. However, the CFO and PN would result in ICI, which increases the difficulty in the estimation of the CFO, PN and the position of the UE. The estimation can only be conducted in the time domain. In this section, we propose an AO-based iterative algorithm, which employs the MAP criterion and GD algorithm to jointly estimate the normalized CFO, PN, and the UE position.
Define the parameter vector $\bm{\lambda}\triangleq\left[\phi,\bm{\theta}^{\top},d_{Ru},\phi_{az},\phi_{el}\right]^{\top}$, wherein $\bm{\theta}$ is a random variable and the others are deterministic. The posterior distribution of $\bm{\lambda}$, given the received signal $\bm{y}$, can be expressed as
@@ III Proposed Joint CFO, PN and UE Position Estimation Algorithm
According to the MAP criterion, $\bm{\lambda}$ can be estimated by optimizing the following unconstrained function:
$\hat{\bm{\lambda}}\propto\arg\min_{\bm{\lambda}}\mathcal{L}(\bm{\lambda}),$ (8)
where $\mathcal{L}(\bm{\lambda})=-\log p(\bm{y}\mid\bm{\lambda})-\log p(\bm{\theta})$, which is obtained by calculating the negative logarithmic function of (7) and then suppressing constants.
@@ III-A1 Approximation of PN vector
Figure 2: Eigenvalues of $\bm{\Psi}$ at the number of subcarriers $N=32$, and the innovation variance for PN $\sigma_{\Delta}^{2}=10^{-3}$.
From the definition of $\bm{\lambda}$, many variables need to be jointly estimated at the UE. To lower the computational complexity of the estimation, we exploit the redundancy of the covariance matrix of the PN vector to approximate it and reduce variables requiring estimation.
As illustrated in Fig. 2, the eigenvalues of $\bm{\Psi}$ are mostly close to zero, indicating significant redundancy in $\bm{\Psi}$. With this property, we can approximate the PN vector as
@@
where $\bm{\eta}\sim\mathcal{N}(\bm{0}_{L\times 1},\mathbf{I}_{L})$ contains $L\leq N$ PN variables, and $\bm{\Pi}\in\mathbb{R}^{N\times L}$ is the transition matrix. We set $L=\frac{N}{2}$ in this paper. To derive $\bm{\Pi}$, we first perform eigenvalue decomposition (EVD) on $\bm{\Psi}$, i.e., $\bm{\Psi}=\mathbf{U}\mathbf{D}\mathbf{U}^{\top}$, where $\mathbf{D}={\rm Diag}(\bm{e})$ and $\bm{e}$ collects the eigenvalues of $\bm{\Psi}$ arranged in descending order. Subsequently, $\bm{\Pi}$ can be calculated by $\bm{\Pi}=[\mathbf{U}]_{:,0:L-1}\times\sqrt{\bm{e}}$. In the following analysis, $\bm{\eta}$ is estimated, instead of $\bm{\theta}$.
@@ III-A2 Estimation of approximate PN vector
By approximating the PN vector $\bm{\theta}$, the objective function on the right-hand side (RHS) of (8) can be rewritten as
@@
where $\bm{\mu}\triangleq\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}}\mathbf{S}^{\top}\bm{h}$ stands for the noise-free received signal at the UE.
Owing to the typically small PN variance of practical oscillators, the elements of the PN matrix $\bm{\Lambda}_{\bm{\theta}}$ can be approximated by employing its first-order Taylor expansion, resulting in $e^{j\bm{\theta}[n]}\approx 1+j\bm{\theta}[n]$. Consequently, we can approximate $\bm{\Lambda}_{\bm{\theta}}$ as $\bm{\Lambda}_{\bm{\theta}}\approx\mathbf{I}_{N}+j{\rm Diag}(\bm{\Pi\eta})$. With this approximation, the objective function in (III-A2) can be rewritten as
@@
where $\bar{\bm{y}}\triangleq\bm{y}-\sqrt{P}\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$, and $\mathbf{Q}\triangleq j\sqrt{P}{\rm Diag}(\hat{\bm{\Lambda}}_{\phi}^{[q]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]})\bm{\Pi}$. The superscript ``$^{q}"$ indicates the $q$-th iteration.
Set the partial derivative of (III-A2) regarding $\bm{\eta}$ to zero, i.e.,
@@ III-B CFO Estimation
To estimate the CFO, we solve the following unconstrained problem:
@@
To enhance the tractability of (14) and derive a closed-form solution, the first-order Taylor approximation is employed. The $n$-th term on the diagonal of $\bm{\Lambda}_{\phi}$ can be approximated as
@@
where $\bm{d}\triangleq\sqrt{P}\hat{\bm{\Lambda}}_{\bm{\theta}}^{[q+1]}\mathbf{S}^{\top}\hat{\bm{h}}^{[q]}$.
Upon setting the partial derivative of (17) with regard to $\phi$ to zero and solving it for $\phi$, we acquire the closed-form solution of $\hat{\phi}$ at the $(q+1)$-th iteration as
@@ III-C Position Estimation
For position estimation, we employ the GD algorithm. Due to the same structures in the respective derivatives of the objective function with regard to the three position parameters $d_{Ru}$, $\phi_{az}$, and $\phi_{el}$, we use the symbol $\xi$ to represent any one of them. The partial derivative of the objective function with regard to $\xi$ is given by
@@
where $\frac{\partial\bm{h}}{\partial\xi}$ is calculated by
@@
$\left.\quad+(d_{Ru}\cos\phi_{el}-z_{r})(-d_{Ru}\sin\phi_{el})\right].$ (23)
III-D Overall Algorithm
The proposed joint estimation of the CFO, PN and UE position is summarized in Algorithm 1. Specifically, we initialize the parameters, including the initial estimation of the normalized CFO $\hat{\phi}$ and UE position $\hat{\bm{p}}_{u}$, the step size, $k_{p}$, the convergence accuracy for outer and inner loops, $\epsilon_{o}$ and $\epsilon_{i}$, and the maximum iteration number for the outer and inner loops, $N_{o}$ and $N_{i}$. After transforming the coordinates into the polar coordinates, we estimate the PN, CFO, and UE position alternately. Closed-form expressions are utilized to update the PN and CFO estimates in each iteration, while the GD algorithm is employed to estimate the UE position. This repeats until either the objective function $\mathcal{L}$ converges or the maximum number of iterations is reached. It is noteworthy that this AO-based algorithm takes the RIS phase shifts, i.e., $\bm{w}$, as input, which can affect the algorithm performance significantly. The configuration of the RIS phase shifts, $\bm{w}$, will be discussed in Section V.
In Algorithm 1, the objective function is expected to be non-increasing throughout iterations when updating the CFO and PN parameters because the approximate functions in (III-A2) and (17) show convexity with respect to $\bm{\eta}$ and $\phi$, respectively. In Steps 9, 10 and 11, with the UE position estimated in the previous iteration as input, the GD algorithm can ensure a lower objective function. Since the posterior probability is no larger than $1$, the negative logarithmic function $\mathcal{L}(\bm{\lambda})$ is non-negative and therefore lower bounded. Algorithm 1 converges.
The computational complexity for updating $\bm{\eta}$, $\bm{\theta}$, $\phi$, and position parameters $d_{Rt}$, $\phi_{az}$, and $\phi_{el}$ in Algorithm 1 are $\mathcal{O}(L^{3}+N^{2}+NL^{2})$, $\mathcal{O}(NL)$, $\mathcal{O}(N^{2})$, and $\mathcal{O}(\max\left\{N^{3},N_{R}\right\})$, respectively. Suppose the inner GD algorithm for updating the UE position converges after $K_{i}$ iterations. Consequently, the computational complexity of Algorithm 1 is $\mathcal{O}(\max\left\{L^{3}+N^{2}+NL^{2},K_{i}N^{3},K_{i}N_{R}\right\})$ per iteration.
Algorithm 1 Proposed iterative joint estimation algorithm.
1:  Input: $\bm{w}$;
2:  Initialization: $\hat{\phi}$, $\hat{\bm{p}}_{u}$, $k_{p}$, $\epsilon_{o}$, $\epsilon_{i}$, $N_{o}$, $N_{i}$;
3:  Calculate polar coordinate of UE using (1);
4:  repeat
5:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\bm{\eta}}$ with $\hat{\phi}$ using (13);
6:     Update $\hat{\bm{\theta}}$ using (9);
7:     Given fixed $\hat{\bm{p}}_{u}$, update $\hat{\phi}$ with $\hat{\bm{\theta}}$ using (18);
8:     repeat
9:        $\hat{d}_{Ru}\leftarrow\hat{d}_{Ru}-\frac{\partial\mathcal{L}_{\xi}}{\partial d_{Ru}}k_{p}$;
10:        $\hat{\phi}_{az}\leftarrow\hat{\phi}_{az}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{az}}k_{p}$;
11:        $\hat{\phi}_{el}\leftarrow\hat{\phi}_{el}-\frac{\partial\mathcal{L}_{\xi}}{\partial\phi_{el}}k_{p}$;
12:     until $\mathcal{L}^{[p]}-\mathcal{L}^{[p+1]}\leq\epsilon_{i}$ after the $p$-th iteration or $p\textgreater N_{i}$.
13:  until $\mathcal{L}^{[q]}-\mathcal{L}^{[q+1]}\leq\epsilon_{o}$ after the $q$-th iteration or $q\textgreater N_{o}$.
14:  Calculate Cartesian coordinate of UE using (2);
15:  Output: $\hat{\phi}$, $\hat{\bm{\theta}}$, $\hat{\bm{p}}_{u}$.
III-E Performance Metrics
We specify the performance metrics for all estimation parameters in $\bm{\lambda}$ for the sake of evaluating Algorithm 1. For the position estimation, we take root mean square error (RMSE) as the performance index. For any testing UE position $\bm{p}_{u}$, ${\rm RMSE}_{u}$ can be expressed as
@@ III-E Performance Metrics
where $M_{e}$ is the number of Monte Carlo tests, and $\hat{\bm{p}}_{u,m}$ denotes the estimation of the UE position in the $m$-th trial.
For the CFO and PN estimation, it can be straightforwardly observed from (6) that the received signal $\bm{y}$ does not change under a common phase rotation $\epsilon$, i.e.,
$\hat{\phi}\rightarrow\phi-\epsilon\quad\text{and }\quad\hat{\bm{\theta}}\rightarrow\bm{\theta}+\bm{\epsilon},$ (25)
where $\bm{\epsilon}[n]=\frac{2\pi n\epsilon}{N}$. Here, (25) indicates that there exists ambiguity in estimating the CFO and PN. The ambiguity leads to difficulties in achieving high estimation accuracy for individual parameters of the CFO and PN. Nevertheless, a high-precision joint estimation of CFO and PN is achievable, which is beneficial for estimating the UE position and will be illustrated in section VI.
To evaluate the joint estimation performance, we utilize a joint MSE measure, which can be calculated as
@@
where $\bm{\underline{\gamma}}=\bm{\gamma}-\bm{\gamma}[0]\bm{1}$, $\hat{\bm{\underline{\gamma}}}=\hat{\bm{\gamma}}-\hat{\bm{\gamma}}[0]\bm{1}$, $\bm{\gamma}=[\bm{\gamma}[0],\bm{\gamma}[1],\cdots,\bm{\gamma}[N-1]]^{\top}$, $\bm{\gamma}[n]=\bm{\theta}[n]+\frac{2\pi n\phi}{N}$, $\hat{\bm{\gamma}}=[\hat{\bm{\gamma}}[0],\hat{\bm{\gamma}}[1],\cdots,\hat{\bm{\gamma}}[N-1]]^{\top}$, and $\hat{\bm{\gamma}}[n]=\hat{\bm{\theta}}[n]+\frac{2\pi n\hat{\phi}}{N}$. Using this measure, the performance of the joint CFO-PN estimation rather than each of the individual variables alone is considered.